OpenStudy (anonymous):

\[\int\limits_{c}^{d} \int\limits_{a}^{b} \frac{\partial^2f}{\partial x\partial y} dxdy = \int\limits_{c}^{d} \left ( \int\limits_{a}^{b} \frac{\partial }{\partial x} \left ( \frac{\partial f}{\partial y} \right ) dx \right ) dy\] ..

OpenStudy (anonymous):

\[\int\limits_{c}^{d}f'(b,y)-f'(a,y)dy\] =f(b,d)-f(b,c)-f(a,d)+f(a,c)